40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 12 16:57:38 2011, Jenne, Update, RF System, RAMmon, 4 day trend EOM_tempSense_noFoam_noHeater_RAMmon_4days.png
    Reply  Tue Dec 13 11:34:32 2011, Jenne, Update, RF System, RAMmon, 5 day trend EOM_tempSense_RAMmon_5days_last_day_withFoam_noHeater.pngEOM_tempSense_withFoam_noHeater_RAMmon_1day.png
Message ID: 6109     Entry time: Mon Dec 12 16:57:38 2011     Reply to this: 6111
Author: Jenne 
Type: Update 
Category: RF System 
Subject: RAMmon, 4 day trend 

EOM was aligned to minimize the 11 and 55 MHz peaks in the RAMmon PD the other day (elog 6089), and was left with just the temperature sensor attached, no heater, no foam box.

Here is a 4 day trend:

EOM_tempSense_noFoam_noHeater_RAMmon_4days.png

I don't have a whole lot to say about this, other than there's a lot of stuff going on.  The craziness at the end is me realigning the PMC and MC since, as you can see, MC trans was way down.  The foam box was put on earlier today (elog 6104), so we'll see how that changes things over night.

ELOG V3.1.3-